OpenStudy (anonymous):

Solve the following logarithmic equation. 2log(base6)(x-5)+log(base6)(9)=2 x=__________________ OR There is no solution.

OpenStudy (anonymous):

\[ 2\log_6(x-5)+\log_6(9)=2 \\ \log_6((x-5)^2)+\log_6(9)=2 \\ \log_6(9(x-5)^2)=2 \\ 6^{\log_6(9(x-5)^2)} =6^2 \\ 9(x-5)^2=36 \]

OpenStudy (anonymous):

thank you super much!